Evaluate the following problems.

1530:Y 11111101 1531 01010101 1532 10000 1111 1533 1000 0000 1534 11111111 1535 10001 1101 1536 0010 1010 1537 1000 0011

a)

1. Evaluate den(Y,0,16)

2. Evaluate hex(Y[0],2)

3. Evaluate HEX(*(Y+4),2)

4. Evaluate oct(*(Y+4),3)

5. Convert Y into 13-bit binary

6. Evaluate den(Y[4],0,2)

7. What is the octal representation (3 digit) of denary 256?

8) What is the Hex ...... (3 Hex digit) of denary 256?
Answer #1

1. Evaluate den(Y,0,16)
1111 1101 0101 0101
0xFD55 = 15*(16^3)+13*(16^2)+5*(16^1)+5*(16^0)
       = 64853

2. Evaluate hex(Y[0],2)
1111 1101 = 0xFD

3. Evaluate HEX(*(Y+4),2)
(value at address (1530+4 = 1534))= 1111 1111
   = 0xFF

4. Evaluate oct(*(Y+4),3)
(Here, we need to take 3 digits at a time, starting from right)
Value at Y[4] = value at 1534 = 1111 1111 = 11 111 111
   = 377 octal

5. Convert Y into 13-bit binary
(Here we need to club 13 digits together)
Y=1111110101010101000011111000000011111111000111010010101010000011
If you convert to 13-digit at a time starting from LSB.
Total digits = 64, so you need to make it multiple of 13 so 65.
One digit is less (65-64). So add a 0 in the beginning and then cut it into set of 13 consecutive digits.
So, Y is as follows:
0111111010101
0101000011111
0000000111111
1100011101001
0101010000011


6. Evaluate den(Y[4],0,2)
Y[4] = 1111 1111 = 0xFF = 15*16 + 15 = 255 decimal

7. What is the octal representation (3 digit) of denary 256?
256 decimal = 100000000 binary
To convert into octal, take 3-digit set from LSB,
100 000 000 binary = 400 octal.

8) What is the Hex ...... (3 Hex digit) of denary 256?
256 decimal = 1 0000 0000 (for hex, take 4-digit set from LSB)
So, 0x100